# Magnetic Polarisability of Octet Baryons via Lattice QCD

### v2

@@ VI.2 Fit window dependence
With confidence in hand that we are choosing fit windows with reasonable suppression of excited state contamination, we still need to select a fit window. This time we consider the proton at $\kappa=0.13727$, $m_{\pi}=570\,$MeV.
The start point criteria derived from the heat maps as described in the previous section requires $t_{s}=29$ in this case. Based purely on the $\chi^{2}_{\rm dof}$ values for the two field strengths, any fit window between $t=29$ and $t=35$ is reasonable. In many cases, such a flexibility in the choice of window produces minimal variation in the resulting value. However, this is not the case here. A summary of the polarisability obtained from various acceptable fit windows are summarised in Tab. 4. Here we see some variation in the resulting magnetic polarisability. This presents a problem as we desire a systematic and consistent scheme through which we can extract the magnetic polarisability. To resolve this issue, we employ a weighted averaging method Beane et al. (2021) to systematically extract a consistent result from a given set of fitting regions.
Table 4: Dependence of the magnetic polarisability of the proton, $\beta_{p}$, on the choice of fit window at $m_{\pi}=570\,$MeV. Magnetic polarisability values in are units of ×$\times 10^{-4}\,$fm3.
$t_{s}$ $t_{f}$ $\beta_{p}$ $\chi^{2}_{{\rm dof},k_{d}=1}$ $\chi^{2}_{{\rm dof},k_{d}=2}$
@@ VI.3 Weighted Averaging
We consider a specific fit window $[t_{s},t_{f}]$ to be eligible if the window has minimum length three ($t_{f}\geq t_{s}+2$) and $t_{f}=t_{\rm max}$. Fixing $t_{f}=t_{\rm max}$ ensures the collection of possible fit windows is sampled equitably. Consideration of every possible end point for a fit starting at a given time slice would favour fit windows commencing at early time slices; they would contribute on more occasions to the weighted average than a fit that starts later. Thus the criteria $t_{f}=t_{\rm max}$ ensures each starting time slice is considered once.
For example in a fitting range $t\in[20,24]$ where all starting points are equally favourable, the possible windows are $[20,22],[20,23],[20,24],[21,23],[21,24],[22,24]$. We clearly see that windows with $t_{s}=20$ would be favoured in the calculation due simply to occurring earlier. Fixing $t_{f}=t_{\rm max}$ provides $[20,24],[21,24]$ and $[22,24]$ which equitably samples possible $t_{s}$ values. This is important as the value of a fit is mostly determined by the first few data points contained in a fit where the uncertainties are small. Over counting early fits is likely to lead to systematic errors.
@@
Here, $\delta E_{i}$ is the uncertainty of the $i$th fit, and $p_{i}$ is the $p$-value of the fit. The $p$-value is the probability of a given fit occurring in a $\chi^{2}$ distribution. It is most easily calculated using the gamma distribution and gamma function in the following manner
$p_{i}=\frac{\Gamma(N_{\rm dof}/2,\chi^{2}/2)}{\Gamma(N_{\rm dof}/2)}\,.$ ((73))
Dividing the degrees of freedom and $\chi^{2}$ of the fit by two and normalising the gamma distribution in the numerator with the gamma function in the denominator causes the gamma distribution to rescale precisely to the $\chi^{2}$ distribution, hence producing the probability of any given fit occurring based on its $\chi^{2}$ and number of degrees of freedom.
@@
We now apply this method to the proton correlators considered in Tab. 4. We assign a start point $t_{\rm min}=29$ based on the heat map of the underlying correlators and choose $t_{\rm max}=34$. Beyond that, all signal has been lost to noise.
The weights for each window are shown in Tab. 5. We can see that the later windows, which are more susceptible to noise due to their larger uncertainties are suppressed.
Table 5: Fit weights of Eq. (71) for the proton at $m_{\pi}=570\,$MeV. Strength is distributed among the candidate fit windows with later fit windows suppressed. Magnetic polarisability values are in units of ×$\times 10^{-4}\,$fm3.
$t_{s}$ $t_{f}$ $\beta_{p}$ $\chi^{2}_{{\rm dof},k_{d}=1}$ $\chi^{2}_{{\rm dof},k_{d}=2}$ $w_{k_{d}=1}$ $w_{k_{d}=2}$
@@
While our procedure assigns the same $t_{\rm min}$ and $t_{\rm max}$ to both $k_{d}$ values governing the field strength, we fit and weight the two field strengths independently. It is not uncommon for their correlators to have significantly different $\chi^{2}_{\rm dof}$ values in spite of their correlated nature. We also explored the possibility of forcing common weights for the field strengths to maximise the opportunity for cancellation of correlated fluctuations. However, we observed this approach to have minimal effect on the final uncertainty. As a result, we fit and weight the two field strengths independently.
@@
The magnetic polarisability is determined in a single parameter fit of
@@
To determine the uncertainty in the polarisability, a jackknife error estimate is performed. A second-order jackknife is used to obtain uncertainties on the correlation functions and correlation function ratios. Uncertainties for fit values such as the magnetic polarisability energy shift for each field strength are obtained from individual first-order jackknife sub-ensembles. Finally, the fit of the energy shifts as a function of field strength is repeated on each jackknife sub-ensemble and the error on the ensemble average calculated as the jackknife error.
@@ VII Lattice Results
Magnetic polarisability values are presented in Tab. 6. and illustrated in Fig. 7 in the context of the constituent quark model of Sec. II. One observes a discrepancy in magnitude between the quark model predictions and the lattice results and we will address this below.
However, a few observations of the lattice results are worthy of note. The uncertainties in the lattice results are sufficiently small to reveal interesting structure in the octet baryon polarisabilities. The $\Xi^{0}$ stands out in magnitude in accord with quark model expectations. This is followed by a cluster of $n$, $p$ and $\Sigma^{+}$ magnetic polarisabilities of similar magnitude, again in accord with the quark model. Finally, the negatively charged $\Xi^{-}$ and $\Sigma^{-}$ baryons have very small polarisabilities as the quark model predicts.
Table 6: Magnetic polarisability results for the outer octet baryons on the PACS-CS ensembles. Polarisability values are in units of ×$\times 10^{-4}\,$fm3, and pion masses are in the Sommer scheme.
@@
Figure 7: Uncorrected (a) and corrected (b) quark model predictions for the magnetic polarisability on the PACS-CS ensembles are displayed as interpolated curves. The lattice data from Tab. 6 are plotted as points on both figures. The legend is common to both plots and ordered to match the vertical ordering at the first dashed line. Dashed lines represent the pion masses of the PACS-CS ensembles used in this work.
VII.1 Corrected Quark Model
To address the magnitude difference between our lattice QCD results and the quark model, we consider a simple correction to the quark model. Recalling Eq. II.1
@@ VII.1 Corrected Quark Model
there are two contributions to the magnetic polarisability. To retain the negative contribution of $\beta_{2}$, we include two positive fit parameters $a_{1},\,a_{2}$
@@
and correct the model by fitting these two parameters to describe the magnetic polarisabilities of all six octet baryons at each of the four quark masses considered in lattice QCD. A simple least squares minimisation produces the fit parameters
$a_{1}=0.417,\quad a_{2}=0.540\,.$ ((79))
We see an approximate reduction in both terms of a half. This corrected model is illustrated in Fig. 7 beside the original uncorrected model. We see that the quark model now broadly agrees in a qualitative manner with the lattice values and provides a detailed understanding of the physics underpinning the magnetic polarisabilities of octet baryons. Given the simplicity of the constituent quark model, the quality of, and insight provided by its predictions are significant.
We see excellent agreement for the $\Xi^{0}\,$. We also see the clear difference in polarisability of the $\Xi^{-}\,$and $\Sigma^{-}\,$as predicted by the quark model, though we lack sufficient precision to make a statement about the sign of those magnetic polarisabilities.
The ordering of baryons in the lattice data is very interesting. Correlated differences between the proton, neutron and $\Sigma^{+}\,$cannot order them uniquely at one sigma. The ordering changes for different quark masses.
We also highlight the quark mass dependence of the proton, neutron and $\Sigma^{+}\,$in the lattice results. We see a clear trend downwards as we approach the physical point. However, this trend is not captured well by the quark model.
@@ VIII.1 Formalism
Figure 8: Two-photon interactions relevant to the baryon magnetic polarisability. In coupling the photons to the intermediate meson, these diagrams generate the most important leading-nonanalytic behaviour in the quark mass dependence of the magnetic polarisability of a baryon. Intermediate baryon states include those degenerate with the initial baryon (left) and those where the intermediate state has a mass different from the initial baryon (right).
To complete our analysis, we now connect our lattice QCD results to the physical world via chiral extrapolation. In doing so, we draw on chiral effective field theory ($\chi$EFT) to inform corrections for systematic uncertainties associated with the finite-volume of the lattice and electro-quenching of the quark sea. This analysis follows the formalism established in Refs. Hall et al. (2014) and Bignell et al. (2020), and extends the formalism to address all six of the outer octet baryons considered herein.
As our lattice QCD results can be described well by a fit linear in $m_{\pi}^{2}$, we consider the following chiral expansion for the magnetic polarisability
@@
where $a_{0}^{\Lambda}$, and $a_{2}^{\Lambda}$ are residual series coefficients which we constrain using volume-corrected lattice QCD results. $\Lambda$ is the renormalisation scale and $\beta^{MB}$ and $\beta^{MB^{\prime}}$ are the leading-order loop contributions to the magnetic polarisability from the diagrams of Fig. 8. The scale-dependent residual-series coefficients are combined with the analytic scale-dependent terms of the loop contributions to recover the renormalised series expansion Young et al. (2003); Hall et al. (2010, 2012, 2013).
@@
where $\omega_{\vec{k}}=\sqrt{\vec{k}^{2}+m_{M}^{2}}$ is the energy carried by the meson $M$ with three-momentum $\vec{k}$, $f_{\pi}=92.4\,$MeV is the pion decay constant and
@@
is a dipole regulator which ensures only soft momenta flow through the effective-field theory degrees of freedom. Eq. (81) generates the leading nonanalytic contribution to the magnetic polarisability proportional to $1/m_{\pi}$. It is through these known contributions that we can correct for the electro-quenched nature of the calculation and estimate finite-volume corrections to the lattice results.
@@
For each of the baryons in this work, $p$, $n$, $\Sigma^{\pm}$, $\Xi^{0,-}$, we consider transitions to all possible octet and decuplet baryons. These transitions involve the mesons $\pi^{\pm,0}$, $K^{\pm,0}$, $\overline{K}^{0}$, $\eta$, and $\eta^{\prime}$. We note that all transitions featuring a neutral meson in the intermediate state vanish in the leading order terms of full QCD, but not in the electro-quenched theory where sea-quark charges are effectively zero.
@@
To correct for the electro-quenching present in the lattice results we must first determine the contributions $\chi_{MB}$ and $\chi_{MB^{\prime}}$ associated with $\beta^{MB^{\prime}}$ and $\beta^{MB^{\prime}}$ respectively. These contributions are derived from the interactions in Fig. 8 with reference to the quark flow diagrams of Fig. 9. There are two cases to examine; one where the quark flow is fully connected and one with a disconnected sea-quark-loop contribution.
@@
Figure 9: The fully connected quark-flow diagram (a) features only valence quarks. The quark-flow disconnected diagram (b) contains a sea-quark-loop contribution to a baryon transition to an intermediate meson-baryon state. In this case, we have a proton, $p$, transitioning to an intermediate $\pi^{+}n$ intermediate state and it is straight forward to extend these quark flow diagrams to consider any of the outer octet baryons with doubly and singly represented quark flavours.
@@
In each case we must attach two photons to the meson that is formed in the interaction. In the connected case, all quarks must be connected to valence quarks, however in the disconnected case we have the option of connecting one or two photons to a disconnected sea quark. As such, we have three types of interactions. We call these valence-valence, valence-sea and sea-sea. In equations, these are abbreviated to their initials.
@@
In the lattice results, only the valence-valence contributions are present. The sea quarks are electrically neutral and diagrams where a photon couples to a sea quark vanish. As such, some contributions to the magnetic polarisability are not present and some contributions which should cancel out in summing the contributions of connected and disconnected quark flows do not do so.
@@
As the chiral coefficients are analytically known, we can proceed by fitting the lattice QCD results in the electro-quenched effective field theory and then explicitly add the missing disconnected sea-quark contributions. Each contribution contributes proportionally with the charge of the interacting quarks, so in the example $p\rightarrow n\,\pi^{+}$ shown in Fig. 9 (right) we have
@@
The coupling strengths may be obtained in a partially-quenched chiral perturbation theory scheme Leinweber (2004) where the flavour of the disconnected sea quark is relabelled to the missing SU(3) quark flavour. In this way the quark disconnected contribution can be isolated in terms of known meson-baryon dressing coefficients. For the case under consideration here, that is the strange quark. In this case the intermediate state would become $\Lambda/\Sigma^{0}$ + $K^{+}$, hence the sea-sea contribution becomes
@@
With the disconnected quark-flow contributions determined, the valence-valence contribution is obtained by subtracting the valence-sea and sea-sea contributions from the full QCD contribution. In some cases the full QCD contribution may be zero leading to equal and opposite contributions from the connected and disconnected quark flow contributions. However, in the case under consideration here, the $p\to n\pi^{+}$ transition makes a contribution proportional to $\chi_{\pi^{+}n}^{2}$ such that
@@
This process is repeated for all possible transitions for each of the outer octet baryons. All relevant chiral coefficients and associated contributions are summarised in Appendix A.
We note that we have not included transitions to an intermediate $\bar{s}s$ pseudoscalar meson. While it does not contribute to full QCD processes, its consideration can in principle lead to a contribution in the process of separating valence and sea contributions. However, the mass of the $\bar{s}s$ pseudoscalar meson is large at approximately $\sqrt{2m_{K}^{2}-m_{\pi}^{2}}\simeq 685\,$MeV. In the finite-range regularisation used here, the contributions from such large-mass mesons are naturally suppressed, such that the partially-quenched correction is negligible.
It’s interesting to note that transitions involving $\pi^{0}$, $\eta$ and $\eta^{\prime}$ intermediate mesons do not contribute to this analysis. In full QCD these mesons are electrically neutral and do not generate a contribution to the leading loop integrals. In partially quenched QCD, their composition of $\bar{u}u$, $\bar{d}d$, and $\bar{s}s$ matched flavours means they are not relevant to the disconnected flow calculations which by definition include a third quark flavour that differs from the quark flavours present in the outer octet baryons under consideration. As such, subtleties associated with SU(3) flavour symmetry breaking in the $\pi^{0}$, $\eta$ and $\eta^{\prime}$ masses do not enter Leinweber and Thomas (2024).

### v3

@@ VI.2 Fit window dependence
With confidence in hand that we are choosing fit windows with reasonable suppression of excited-state contamination, we still need to select a fit window. This time we consider the proton at $\kappa=0.13727$, $m_{\pi}=570\,$MeV.
The start point criteria derived from the heat maps as described in the previous section requires $t_{s}=29$ in this case. Based purely on the $\chi^{2}_{\rm d.o.f.}$ values for the two field strengths, any fit window between $t=29$ and $t=35$ is reasonable. In many cases, such a flexibility in the choice of window produces minimal variation in the resulting value. However, this is not the case here. A summary of the polarizability obtained from various acceptable fit windows are summarized in Table 4. Here we see some variation in the resulting magnetic polarizability. This presents a problem as we desire a systematic and consistent scheme through which we can extract the magnetic polarizability. To resolve this issue, we employ a weighted averaging method Beane et al. (2021) to systematically extract a consistent result from a given set of fitting regions.
Table 4: Dependence of the magnetic polarizability of the proton $\beta_{p}$ on the choice of fit window at $m_{\pi}=570\,$MeV. Magnetic polarizability values in are units of ×$\times 10^{-4}\,$fm3.
$t_{s}$ $t_{f}$ $\beta_{p}$ $\chi^{2}_{{\rm d.o.f.},k_{d}=1}$ $\chi^{2}_{{\rm d.o.f.},k_{d}=2}$
@@ VI.3 Weighted averaging
We consider a specific fit window $[t_{s},t_{f}]$ to be eligible if the window has minimum length three ($t_{f}\geq t_{s}+2$) and $t_{f}=t_{\rm max}$. Fixing $t_{f}=t_{\rm max}$ ensures the collection of possible fit windows is sampled equitably. Consideration of every possible end point for a fit starting at a given time slice would favour fit windows commencing at early time slices; they would contribute on more occasions to the weighted average than a fit that starts later. The criteria $t_{f}=t_{\rm max}$ ensures each starting time slice is considered once.
For example in a fitting range $t\in[20,24]$ where all starting points are equally favorable, the possible windows are $[20,22],[20,23],[20,24],[21,23],[21,24],[22,24]$. We clearly see that windows with $t_{s}=20$ would be favored in the calculation due simply to occurring earlier. Fixing $t_{f}=t_{\rm max}$ provides $[20,24],[21,24]$ and $[22,24]$ which equitably samples possible $t_{s}$ values. This is important as the value of a fit is mostly determined by the first few data points contained in a fit where the uncertainties are small. Over counting early fits is likely to lead to systematic errors.
@@
Here, $\delta E_{i}$ is the uncertainty of the $i$th fit, and $p_{i}$ is the $p$-value of the fit. The $p$-value is the probability of a given fit occurring in a $\chi^{2}$ distribution. It is most easily calculated using the $\gamma$ distribution and $\gamma$ function in the following manner
$p_{i}=\frac{\Gamma(N_{\rm d.o.f.}/2,\chi^{2}/2)}{\Gamma(N_{\rm d.o.f.}/2)}\,.$ ((73))
Dividing the degrees of freedom and $\chi^{2}$ of the fit by two and normalising the $\gamma$ distribution in the numerator with the $\gamma$ function in the denominator causes the gamma distribution to rescale precisely to the $\chi^{2}$ distribution, hence producing the probability of any given fit occurring based on its $\chi^{2}$ and number of degrees of freedom.
@@
We now apply this method to the proton correlators considered in Table 4. We assign a start point $t_{\rm min}=29$ based on the heat map of the underlying correlators and choose $t_{\rm max}=34$. Beyond that, all signal has been lost to noise.
The weights for each window are shown in Table 5. We can see that the later windows, which are more susceptible to noise due to their larger uncertainties are suppressed.
Table 5: Fit weights of Eq. (71) for the proton at $m_{\pi}=570\,$MeV. Weight is distributed among the candidate fit windows with later fit windows suppressed. Magnetic polarizability values are in units of ×$\times 10^{-4}\,$fm3.
$t_{s}$ $t_{f}$ $\beta_{p}$ $\chi^{2}_{{\rm d.o.f.},k_{d}=1}$ $\chi^{2}_{{\rm d.o.f.},k_{d}=2}$ $w_{k_{d}=1}$ $w_{k_{d}=2}$
@@
While our procedure assigns the same $t_{\rm min}$ and $t_{\rm max}$ to both $k_{d}$ values governing the field strength, we fit and weight the two field strengths independently. It is not uncommon for their correlators to have significantly different $\chi^{2}_{\rm d.o.f.}$ values in spite of their correlated nature. We also explored the possibility of forcing common weights for the field strengths to maximize the opportunity for cancellation of correlated fluctuations. However, we observed this approach to have minimal effect on the final uncertainty. As a result, we fit and weight the two field strengths independently.
@@
The magnetic polarizability is determined in a single-parameter fit of
@@
To determine the uncertainty in the polarizability, a jackknife error estimate is performed. A second-order jackknife is used to obtain uncertainties on the correlation functions and correlation function ratios. Uncertainties for fit values such as the magnetic polarizability energy shift for each field strength are obtained from individual first-order jackknife subensembles. Finally, the fit of the energy shifts as a function of field strength is repeated on each jackknife sub-ensemble and the error on the ensemble average calculated as the jackknife error.
@@ VII Lattice Results
Magnetic polarizability values are presented in Table 6 and illustrated in Fig. 7 in the context of the constituent quark model of Sec. II. One observes a discrepancy in magnitude between the quark model predictions and the lattice results and we will address this below.
However, a few observations of the lattice results are worthy of note. The uncertainties in the lattice results are sufficiently small to reveal interesting structure in the octet-baryon polarizabilities. The $\Xi^{0}$ stands out in magnitude in accord with quark model expectations. This is followed by a cluster of $n$, $p$, and $\Sigma^{+}$ magnetic polarizabilities of similar magnitude, again in accord with the quark model. Finally, the negatively charged $\Xi^{-}$ and $\Sigma^{-}$ baryons have very small polarizabilities as the quark model predicts.
Table 6: Magnetic polarizability results for the outer octet baryons on the PACS-CS ensembles. Polarizability values are in units of ×$\times 10^{-4}\,$fm3, and pion masses are in the Sommer scheme as described in subsection V.1.
@@
Figure 7: Original (a) and scaled (b) quark model predictions for the magnetic polarizability on the PACS-CS ensembles are displayed as interpolated curves. The lattice data from Table 6 are plotted as points on both figures. The legend is common to both plots and ordered to match the vertical ordering at the first dashed line. Dashed lines represent the pion masses of the PACS-CS ensembles used in this work.
VII.1 Improved quark model
To address the magnitude difference between our lattice QCD results and the quark model, we consider a simple scaling of the quark model. Recalling Eq. II.1
@@ VII.1 Improved quark model
there are two contributions to the magnetic polarizability. To retain the negative contribution of $\beta_{2}$, we include two positive fit parameters $a_{1},\,a_{2}$
@@
and scale the model by fitting these two parameters to describe the magnetic polarizabilities of all six octet baryons at each of the four quark masses considered in lattice QCD. A simple least squares minimization produces the fit parameters
$a_{1}=0.401,\quad a_{2}=0.532\,.$ ((79))
We see an approximate reduction in both terms of a half.
There are several factors beyond the considerations of the simple quark model model presented here that can combine to generate corrections of this magnitude. Consider for example the quark model calculation of Ref. Darewych et al. (1983) for the photon-decay amplitudes and widths of low-lying strange baryons, generalized to the outer members of the baryon octet in Ref. Leinweber et al. (1993). Here additional kinematic factors are taken into account
$\mu_{\mathcal{B}\mathcal{B}^{*}}=\frac{2\sqrt{2}}{3}\sqrt{\frac{m_{\mathcal{B}}}{m_{\mathcal{B}^{*}}}}\,e^{-K}\left(\mu_{D}-\mu_{S}\right)\,.$ ((80))
Comparing with Eq. (23), the mass-ratio and $e^{-K}$ factors are new and combine to generate a suppression of $\sim 15$% in the transition moment at the physical point, a suppression approaching 30% upon squaring in Eq. VII.1 for the transition term, $\beta_{1}$.
Furthermore, the $\Delta$ baryon is a resonance observed in pion-nucleon scattering. Its structure is richer than a simple three-quark single-particle state. Two-particle $\pi N$ components mix with the single-particle component of the quark model to form the resonance and this physics is not considered in the simple model presented here. The lattice results suggest that these components act to further suppress the transition moment.
With regard to the suppression of the quark distribution term $\beta_{2}$, we note that a reduction to 70% of the rms charge radius provides a 50% correction in $\beta_{2}$ as the rms charge radius is squared in Eq. VII.1. In obtaining values for the quark distribution radii, we have referred to results from lattice QCD. As such, the radii will include contributions from the meson cloud that dresses the state and becomes an integral part of its structure. One could argue that the radii to be used in $\beta_{2}$ should be the distribution radii of the constituent quark degrees of freedom within the confinement potential. These radii, void of meson cloud contributions, would be smaller and could account for the suppression factor derived from the lattice QCD polarizabilities.
The scaled model is illustrated in Fig. 7 beside the original quark model. We see that the quark model now broadly agrees in a qualitative manner with the lattice values and provides a detailed understanding of the physics underpinning the magnetic polarizabilities of octet baryons. Given the simplicity of the constituent quark model, the quality of and insight provided by its predictions are significant.
We see excellent agreement for the $\Xi^{0}\,$. We also see the clear difference in polarizability of the $\Xi^{-}\,$and $\Sigma^{-}\,$as predicted by the quark model, though we lack sufficient precision to make a statement about the sign of those magnetic polarizabilities.
The ordering of baryons in the lattice data is very interesting. Correlated differences between the proton, neutron and $\Sigma^{+}\,$cannot order them uniquely at $1\sigma$. The ordering changes for different quark masses.
We also highlight the quark mass dependence of the proton, neutron, and $\Sigma^{+}\,$in the lattice results. We see a clear trend downward as we approach the physical point. However, this trend is not captured well by the quark model.
@@ VIII.1 Formalism
Figure 8: Two-photon interactions relevant to the baryon magnetic polarizability. In coupling the photons to the intermediate meson, these diagrams generate the most important leading nonanalytic behavior in the quark mass dependence of the magnetic polarizability of a baryon. Intermediate baryon states include those degenerate with the initial baryon (left) and those where the intermediate state has a mass different from the initial baryon (right).
To complete our analysis, we now connect our lattice QCD results to the physical world via chiral extrapolation. In doing so, we draw on chiral effective field theory ($\chi$EFT) to inform corrections for systematic uncertainties associated with the finite volume of the lattice and electroquenching of the quark sea. This analysis follows the formalism established in Refs. Hall et al. (2014) and Bignell et al. (2020) and extends the formalism to address all six of the outer octet baryons considered herein.
As our lattice QCD results can be described well by a fit linear in $m_{\pi}^{2}$, we consider the following chiral expansion for the magnetic polarizability
@@
where $a_{0}^{\Lambda}$, and $a_{2}^{\Lambda}$ are residual series coefficients which we constrain using volume-corrected lattice QCD results. $\Lambda$ is the renormalization scale and $\beta^{MB}$ and $\beta^{MB^{\prime}}$ are the leading-order loop contributions to the magnetic polarizability from the diagrams of Fig. 8. The scale-dependent residual series coefficients are combined with the analytic scale-dependent terms of the loop contributions to recover the renormalized series expansion Young et al. (2003); Hall et al. (2010, 2012, 2013).
@@
where $\omega_{\vec{k}}=\sqrt{\vec{k}^{2}+m_{M}^{2}}$ is the energy carried by the meson $M$ with three-momentum $\vec{k}$, $f_{\pi}=92.4\,$MeV is the pion-decay constant, and
@@
is a dipole regulator which ensures only soft momenta flow through the effective field theory degrees of freedom. Equation (82) generates the leading nonanalytic contribution to the magnetic polarizability proportional to $1/m_{\pi}$. It is through these known contributions that we can correct for the electroquenched nature of the calculation and estimate finite-volume corrections to the lattice results.
@@
For each of the baryons in this work, $p$, $n$, $\Sigma^{\pm}$, $\Xi^{0,-}$, we consider transitions to all possible octet and decuplet baryons. These transitions involve the mesons $\pi^{\pm,0}$, $K^{\pm,0}$, $\overline{K}^{0}$, $\eta$, and $\eta^{\prime}$. We note that all transitions featuring a neutral meson in the intermediate state vanish in the leading-order terms of full QCD, but not in the electroquenched theory where sea-quark charges are effectively zero.
@@
To correct for the electroquenching present in the lattice results we must first determine the contributions $\chi_{MB}$ and $\chi_{MB^{\prime}}$ associated with $\beta^{MB^{\prime}}$ and $\beta^{MB^{\prime}}$, respectively. These contributions are derived from the interactions in Fig. 8 with reference to the quark-flow diagrams of Fig. 9. There are two cases to examine; one where the quark flow is fully connected and one with a disconnected sea-quark-loop contribution.
@@
Figure 9: The fully connected quark-flow diagram (a) features only valence quarks. The quark-flow disconnected diagram (b) contains a sea-quark-loop contribution to a baryon transition to an intermediate meson-baryon state. In this case, we have a proton $p$ transitioning to an intermediate $\pi^{+}n$ intermediate state. It is straightforward to extend these quark-flow diagrams to consider any of the outer octet baryons with doubly and singly represented quark flavors.
@@
In each case we must attach two photons to the meson that is formed in the interaction. In the connected case, all quarks must be connected to valence quarks, however in the disconnected case we have the option of connecting one or two photons to a disconnected sea quark. As such, we have three types of interactions. We call these ”valence-valence”, ”valence-sea” and ”sea-sea”. In equations, these are abbreviated to their initials.
@@
In the lattice results, only the valence-valence contributions are present. The sea quarks are electrically neutral and diagrams where a photon couples to a sea quark vanish. As such, some contributions to the magnetic polarizability are not present and some contributions which should cancel out in summing the contributions of connected and disconnected quark flows do not do so.
@@
As the chiral coefficients are analytically known, we can proceed by fitting the lattice QCD results in the electroquenched effective field theory and then explicitly add the missing disconnected sea-quark contributions. Each contribution contributes proportionally with the charge of the interacting quarks, so in the example $p\rightarrow n\,\pi^{+}$ shown in Fig. 9 (right) we have
@@
The coupling strengths may be obtained in a partially quenched chiral perturbation theory scheme Leinweber (2004) where the flavor of the disconnected sea quark is relabeled to the missing SU(3) quark flavor. In this way the quark disconnected contribution can be isolated in terms of known meson-baryon dressing coefficients. For the case under consideration here, that is the strange quark. In this case the intermediate state would become $\Lambda/\Sigma^{0}$ + $K^{+}$, hence the sea-sea contribution becomes
@@
With the disconnected quark-flow contributions determined, the valence-valence contribution is obtained by subtracting the valence-sea and sea-sea contributions from the full QCD contribution. In some cases the full QCD contribution may be zero leading to equal and opposite contributions from the connected and disconnected quark-flow contributions. However, in the case under consideration here, the $p\to n\pi^{+}$ transition makes a contribution proportional to $\chi_{\pi^{+}n}^{2}$ such that
@@
This process is repeated for all possible transitions for each of the outer octet baryons. All relevant chiral coefficients and associated contributions are summarized in Appendix A.
We note that we have not included transitions to an intermediate $\bar{s}s$ pseudoscalar meson. While it does not contribute to full QCD processes, its consideration can in principle lead to a contribution in the process of separating valence and sea contributions. However, the mass of the $\bar{s}s$ pseudoscalar meson is large at approximately $\sqrt{2m_{K}^{2}-m_{\pi}^{2}}\simeq 685\,$MeV. In the finite-range regularization used here, the contributions from such large-mass mesons are naturally suppressed, such that the partially quenched correction is negligible.
It is interesting to note that transitions involving $\pi^{0}$, $\eta$, and $\eta^{\prime}$ intermediate mesons do not contribute to this analysis. In full QCD these mesons are electrically neutral and do not generate a contribution to the leading loop integrals. In partially quenched QCD, their composition of $\bar{u}u$, $\bar{d}d$, and $\bar{s}s$ matched flavors means they are not relevant to the disconnected flow calculations which by definition include a third quark flavor that differs from the quark flavors present in the outer octet baryons under consideration. As such, subtleties associated with SU(3) flavor symmetry breaking in the $\pi^{0}$, $\eta$, and $\eta^{\prime}$ masses do not enter Leinweber and Thomas (2024).